OpenStudy (anonymous):

4-3). Diane has $1.85 in dimes and nickels. She has a total of 24 coins. How many of each kind does she have? how many nickles? how many dimes?

OpenStudy (anonymous):

Let the nickels be x and the dimes be y x + y = 24 5x + 10y = 185 5x + 5y = 120 5y = 65 y=13 x=11 Answer: 11 nickels and 5 dimes